OpenStudy (anonymous):

It takes a machine 15 minutes to put labels on 300 cans of soup. At this rate, How many minutes will it take the machine to put labels on 500 cans of soup?? A. 25 minutes B. 33.3 minutes C. 45 minutes D. 75 minutes

OpenStudy (wach):

Ok, so fist, try and find the rate per hundred cans of soup.

OpenStudy (wach):

If it takes 15 minutes to label 300 cans, that means that in one minute, x amount of cans are labeled, right?

OpenStudy (wach):

Try finding the amount of time it takes to label one can, and then multiply that by the number of cans needed - 500.

OpenStudy (anonymous):

wow!!! thats hard!!!

OpenStudy (anonymous):

i dont know!!!

OpenStudy (wach):

Okay, so how would you try and find the amount of time it takes to label 100 cans? If you know it takes 15 minutes to label 300?

OpenStudy (anonymous):

A.25 minutes. This problem is comparing two rates, when we are comparing rates, we can use proportions. If the machine can put 300 labels in 15 min, then how long would it take to put 500 labels. 300 cans/ 15 min = 500 cans/ x min. They are proportional We can now cross multiply, 300 x= 500*15, we get 300 x= 7500, we now solve for x, divide 7500/ 300, and now x= 25. Final answer, it takes 25 minutes for the machine to put 500 labels.

OpenStudy (anonymous):

i got 25!!!

OpenStudy (anonymous):

Yay! you got it.